Problem 3-10 (LO. 2)

Cortez, Inc., reports $600,000 of pretax book net income in 2019. Cortez’s book depreciation exceeds tax depreciation that year by $20,000. The corporation reports no other temporary or permanent book-tax differences. Cortez’s pertinent U.S. tax rate is 21%, and Cortez earns an after-tax rate of return on capital of 8%.

For 2019, compute the following for Cortez, Inc.

a. Current income tax expense $
b. Deferred income tax expense $
c. Total income tax expense $

a. not 620,000, b. not 130,200, c. is 126,000
Answer #1
Taxable income 620000 =600000+20000
a. Current income tax expense 130200 =620000*21%
b. Deferred income tax expense (4200) =20000*21%
c. Total income tax expense 126000 =600000*21%